# Collaborative Chain-of-Agents for Parametric-Retrieved Knowledge Synergy

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a promising framework for enhancing the capabilities of Large Language Models (LLMs), especially in knowledge-intensive tasks. Despite its advantages, current RAG methods often struggle to fully exploit knowledge during generation. In particular, the synergy between the model’s internal parametric knowledge and external retrieved knowledge remains limited. Retrieved contents may sometimes mislead generation, while certain generated content can guide the model toward more accurate outputs. In this work, we propose Collaborative Chain-of-Agents, a framework designed to enhance explicitly synergy over both parametric and retrieved knowledge. Specifically, we first introduce CoCoA-zero, a multi-agent RAG framework that first performs conditional knowledge induction and then reasons answers. Building on this, we develop CoCoA, a long-chain training strategy that synthesizes extended multi-agent reasoning trajectories from CoCoA-zero to fine-tune the LLM. This strategy enhances the model’s capability to explicitly integrate and jointly leverage parametric and retrieved knowledge. Experiments results show that CoCoA-zero and CoCoA achieve superior performance on open-domain and multi-hop QA tasks.  [^1] Code available at https://github.com/liunian-Jay/CoCoA .
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al. 2023; Touvron et al. 2023) have demonstrated strong performance across a wide range of natural language tasks. However, the knowledge they rely on is embedded in their parameters and cannot be easily updated as new information emerges (Ji et al. 2023; He, Zhang, and Roth 2022). To address this limitation, the Retrieval Augmented Generation (RAG) framework introduces an external retrieval component that brings in external knowledge and integrates it into the input context of the LLMs. This design has led to notable improvements in various natural language processing applications (Gao et al. 2023; Lewis et al. 2020). Existing research has primarily aimed to improve two aspects of RAG: retrieving more relevant information during retrieval and better utilizing that information to guide generation during generation. Despite these efforts, most retrieval-augmented language models (RALMs) still emphasize external retrieval, while paying insufficient attention to the rich internal knowledge already encoded in model parameters. This internal knowledge is especially valuable for open-domain question answering, where many queries are factual and often already covered during pretraining.
Figure 1: Evaluation on 2WikiMultiHopQA, HotpotQA, and WebQuestions. The Merge method is a simple strategy we use to verify the collaboration of internal and external knowledge. It directly generates a passage and merges it into the retrieved passages as the context of the LLM.
Specifically, as the knowledge in LLM’s parameter becomes richer and the abilitiy of the LLM becomes stronger, sometimes answers with search information are not as good as direct answers. To validate the necessity of collaboratively synergizing internal (or parametric) and external (or retrieved) knowledge, we conducted experiments to compare performance. As shown in Fig. 1, across the three evaluation tasks, direct generation and GenRead (Yu et al. 2022) (explicitly generated content) sometimes shows stronger performance. Also, we conduct a test experiment,“Merge”, that explicitly integrates internal and external knowledge by combining retrieved passages with internally generated passages as the final context, as shown in Fig. 1. “Merge” often achieves better results than both direct generation and RAG approaches, demonstrating the potential of internal and external knowledge collaboration. However, its improvements are not consistent across all datasets, indicating the need for more sophisticated integration methods.
@@
Existing methods solve the problem of knowledge collaboration through RAG pipeline optimization. Some approaches alleviate this through workflow or multi-module collaboration. For example, SURE (Kim et al. 2024a) generates multiple candidate answers and verifies them one by one to ensure reliability. CON (Yu et al. 2023) mitigates external noise by adding a processing chain. AstuteRAG (Wang et al. 2024a) integrates reliable information iteratively. There are also some approaches that solve the problem of knowledge collaboration through enhanced training of the LLM. For instance, RAFT (Zhang et al. 2024) employs anti-noise training to enable the model to effectively utilize internal knowledge when external documents contain noise, while Self-RAG (Asai et al. 2023) learns to determine whether retrieval is needed in advance, thereby avoiding harmful content before retrieval. Despite these efforts, existing work still has notable limitations. On the one hand, some pipeline optimization methods tend to lose effectiveness as LLMs become more capable. On the other hand, some fine-tuning methods cannot fully exploit internal knowledge effectively.
@@
In general, our contributions are summarized as follows:
@@ 2.1 Preliminaries
We formalize the standard Retrieval-Augmented Generation framework. Given a query $q$ and a corpus $\mathcal{D}$, the RAG system retrieves $k$ relevant passages $C=\{c_{1},c_{2},\cdots,c_{k}\}\subset\mathcal{D}$ and generates an answer $\hat{a}$ based on the combined input. This process follows a retrieve-then-generate paradigm and can be formulated as:
$\begin{split}C&=\mathcal{R}(q,\mathcal{D},k),\\ \hat{a}&=\mathcal{G}(\mathcal{P}(q,C)),\end{split}$ (1)
where $\mathcal{R}$ is the retriever, $\mathcal{P}$ is the prompt constructor that formats $q$ and $C$, and $\mathcal{G}$ is the generator (e.g., a LLM) that predicts the final answer $\hat{a}$.
2.2 Two-stage RAG Framework: CoCoA-zero
In this section, we present our multi-agent RAG framework, CoCoA-zero, which also functions as the data synthesis pipeline for CoCoA. As shown in Fig. 2, Stage I (§ 2.2) employs two specialized agents to induce knowledge from parameters and retrieval, while Stage II (§ 2.2) introduces a agent to synthesize them for high-level decision-making.
Stage I: Knowledge Induction.
It is challenging to extract implicit knowledge solely from the model’s internal knowledge or retrieved passages. Inspired by GenRead (Yu et al. 2022) and SURE (Kim et al. 2024a), we design two dedicated agents for knowledge induction. Each agent first generates an answer to the question and then summarizes knowledge based on that answer.
Induction of Internal Knowledge. Directly allowing the model to explicitly generate its own internal knowledge is difficult to control and will inevitably result in sparse or inconsistent knowledge being generated. Following SURE (Kim et al. 2024a), we introduce conditional induction. Specifically, the Internal Knowledge Agent samples a candidate $a_{\text{in}}$ from the LLM based on the question:
@@ 2.2 Two-stage RAG Framework: CoCoA-zero
Induction of External Knowledge. For retrieved passages, the External Knowledge Agent follows a similar procedure. Specially, it first retrieve some passages $C=\{c_{1},c_{2},\cdots,c_{k}\}$ from the corpus $\mathcal{D}$. Conditioned on both $q$ and $C$, it produces a second candidate $a_{\text{ext}}$:
@@
The conditional knowledge induction framework makes implicit knowledge explicit and controllable, improving the model’s ability to express knowledge and providing a solid basis for high-level decision-making in the next stage.
Stage II: High-level Decision Making.
@@
Specifically, the Decision-Making Agent adopts COT (Wei et al. 2022) reasoning over the internal and external candidate answers and their corresponding knowledge. It will be prompted with all five components (questions, internal and external candidate answers and their corresponding inductive knowledge) and generate the final answer $\hat{a}$ through COT.
@@ 2.3 Collaborative Chain-of-Agents Training
Figure 3: Illustration of the training for CoCoA.
Supervised Fine-Tuning.
@@
This training explicitly exposes the model to long collaborative samples, where the target outputs are synthesized based on the outputs of CoCoA-zero. Through end-to-end training, multiple agents can influence and enhance each other’s capabilities. Moreover, the noise introduced by intermediate agents becomes negligible, as it contributes to the overall robustness of the training process.
@@
2.4 Optimization Analysis
We compare independent training and CoCoA training under a simplified two-step setting involving pre-generation processing followed by answer generation:
$\mathcal{L}{\text{indep}}=-\log P_{\theta}(s|x,d)-\log P_{\phi}(\hat{a}|s)$ (9)
$\mathcal{L}{\text{chain}}=-\log P_{\theta}(s|x,d)-\log P_{\theta}(\hat{a}|x,d,s)$ (10)
Gradient comparison:
$\frac{\partial\mathcal{L}{\text{chain}}}{\partial\theta}=\frac{\partial\mathcal{L}{\text{indep}}}{\partial\theta}+\Delta_{g}$ (11)
where $\Delta_{g}:=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}|x,s,d)\bigr{]}$. $\Delta_{g}$ captures feedback from the answer to the pre-processing, which is absent in independent training. Chain training is a special type of multi-task learning that helps to break out of local optimization. The experimental results are in Section 3.6, and detailed derivations are in Appendix D.
3 Experiments
In this section, we report our experiments results, and provide a analysis of them.
3.1 Implementation Details
Training Data
We sample subsets from the training sets of HotpotQA (Ho et al. 2020a), 2WikiMultiHopQA (Ho et al. 2020b) and WebQuestions (Berant et al. 2013), then synthesize data using the CoCoA-zero and filter them based on gold answers. This results in 6.8k filtered samples for SFT. For DPO, we select 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero framework. For each sample, we gather 5 relevant passages using CONTRIEVER (Izacard et al. 2021).
Training Details
We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05). During SFT, we train for 5 epochs with a learning rate of 3e-5. For DPO, we used $\beta$=0.2 and $\alpha$=0.2 (RPO), with a learning rate of 5e-6. All experiments are conducted on a single A100 GPU.
Inference Details
During inference, we use Contriever (Izacard et al. 2021) as the retriever and set k to 5. For all datasets, we use 21M English Wikipedia (Karpukhin et al. 2020) dump as the source passages for the retrieval. Prompts for the experiments can be found in Appendix F.
3.2 Datasets and Evaluation Metrics
Eval Datasets To evaluate the effectiveness and generalization of CoCoA, we conduct experiments on two open-domain question answering datasets: WebQuestions (Berant et al. 2013), and TriviaQA (Joshi et al. 2017), as well as two multi-hop question answering benchmarks: HotpotQA (Ho et al. 2020a) and 2WikiMultiHopQA (Ho et al. 2020b). Dataset statistics are summarized in Table 2, and further details are provided in Appendix A.
Task Type Datasets # Samples
Multi-HopQA 2WikiMultiHopQA 500
HotpotQA 500
OpenQA WebQuestions 2032
TriviaQA 11313
Table 2: Description of tasks and evaluation datasets.
Evaluation Metrics
We report both exact match (EM) and F1 scores. Following Asai et al. (2023); Mallen et al. (2022), we adopt a non-strict EM metric that deems a prediction correct if it contains the gold answer. F1 measures token-level overlap between the predicted and gold answers. In our setting, longer responses often yield higher EM scores due to increased coverage, but may reduce F1 scores by introducing irrelevant content. Thus, considering both metrics provides a more balanced evaluation.
@@ 3.3 Baselines
We selected several of the most representative methods for comparison. 1) StandardRAG, which is the most classic “retrieve-then-read” paradigm. 2) Chain-Of-Thought (Wei et al. 2022): Uses CoT prompting to generate reasoning steps before producing the final answer. 3) Chain-Of-Note (Yu et al. 2023): Refines the retrieved passages prior to answering. 4) GenRead (Yu et al. 2022): Generates self-contained intermediate context to answer, effectively replacing retrieval with generation. 5) SURE (Kim et al. 2024a): Conditional summarization followed by multiple validation. 6) Self-RAG (Asai et al. 2023): Employs adaptive retrieval and self-reflection to decide when and how to use external knowledge. 7) DeepSeek-R1-Distill-8B (Guo et al. 2025): A distilled LLaMA-8B model released by DeepSeek-R1, trained on reasoning data. 8) InstructRAG (Wei, Chen, and Meng 2024): Denoising training using self-synthesized data. All retrieval-based methods use top-5 passages. Other experimental settings follow those reported in the original papers. Other experimental settings are shown in the Appendix B.
@@ 3.4 Main Results
Retrieval vs. non-retrieval methods
On WikiMQA and WebQuestions, direct generation performs better, while retrieval methods excel on other tasks. This demonstrates that retrieved knowledge and parametric knowledge each have their own strengths and weaknesses in different scenarios.
RAG without training.
The improvements of some process optimization methods are decreasing compared to standardRAG. We speculate that this is because current LLMs are becoming more powerful enough to make good use of external knowledge. CoCoA-zero improves the average EM and F1 of all tasks by 4.99% and 4.64% respectively, while other train-free methods show little effect. These results suggest that current QA tasks should place greater emphasis on leveraging the model’s rich internal knowledge.
Superiority and Generalization of CoCoA.
Our CoCoA methods achieve state-of-the-art performance across almost all datasets. In particular, CoCoA improves the EM and F1 of 2WikiMultiHopQA tasks by 15.2% and 15.51% respectively. Moreover, despite being trained with limited data, CoCoA also performed well on other out-of-distribution datasets like TriviaQA, demonstrating its robustness.
Reasoning Distillation vs. CoCoA Training.
DeepSeek-R1-8B, trained on distilled reasoning data, outperforms the undistilled StandardRAG. CoCoA, distilled with multi-agent self-synthesis on knowledge-intensive tasks, further surpasses DeepSeek-R1-8B. We speculate this is because logical reasoning and knowledge-intensive tasks differ, and CoCoA can better leverage knowledge. This suggests that explicitly leveraging key internal and external knowledge can be more effective than chain-of-thought reasoning.
Benefit of DPO Training.
Comparing our supervised and DPO variants, DPO training yields improvements across several datasets. This suggests that contrastive preference learning can help the model better align to the collaborative responses of multi-agents. However, it may also lead to performance degradation due to the quality of training data.
Figure 4: Performance varies with the number of documents: 2WikiMQA (left), HotpotQA (middle), WebQuestions (right).
3.5 Ablation Study I: Different Modules

### v3

# CoCoA: Collaborative Chain-of-Agents for Parametric-Retrieved Knowledge Synergy

@@ Abstract
Retrieval-Augmented Generation (RAG) enhances Large Language Models (LLMs), especially for knowledge-intensive tasks. Despite its advantages, current RAG methods often struggle to fully exploit knowledge during generation. In particular, the synergy between the model’s internal parametric knowledge and external retrieved knowledge remains limited. Retrieved contents may sometimes mislead generation, while certain generated content can guide the model toward more accurate outputs. In this work, we propose Collaborative Chain-of-Agents, a framework designed to enhance explicitly synergy over both parametric and retrieved knowledge. Specifically, we first introduce CoCoA-zero, a multi-agent RAG framework that first performs conditional knowledge induction and then reasons answers. Building on this, we develop CoCoA, a long-chain training strategy that synthesizes extended multi-agent reasoning trajectories from CoCoA-zero to fine-tune the LLM. This strategy enhances the model’s capability to explicitly integrate and jointly leverage parametric and retrieved knowledge. Experimental results demonstrate the superiority of CoCoA in open-domain QA and multi-hop QA. Code is public [^1] Code available at https://github.com/liunian-Jay/CoCoA ..
CoCoA: Collaborative Chain-of-Agents for Parametric-Retrieved Knowledge Synergy
Yi Jiang1, Sendong Zhao1thanks: Corresponding author., Jianbo Li1, Haochun Wang1, Lizhe Zhang2, Yan Liu2, Bing Qin1 1Harbin Institute of Technology, China 2China Mobile Group Heilongjiang Co.,Ltd {yjiang,sdzhao,jbli,hcwang,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have demonstrated strong performance across a wide range of natural language tasks. However, the knowledge they rely on is embedded in their parameters and cannot be easily updated as new information emerges (Ji et al., 2023; He et al., 2022). To address this limitation, the Retrieval Augmented Generation (RAG) framework introduces an external retrieval component that brings in external knowledge and integrates it into the input context of the LLMs. This design has led to notable improvements in various natural language processing applications (Gao et al., 2023; Lewis et al., 2020). Existing research has primarily aimed to improve two aspects of RAG: retrieving more useful information during retrieval and better utilizing information to guide generation during generation. Despite these efforts, most retrieval-augmented language models (RALMs) still emphasize external retrieval, while paying insufficient attention to the rich internal knowledge already encoded in model parameters. This internal knowledge is especially valuable for open-domain question answering, where many queries are factual and often already covered during pretraining.
Figure 1: Evaluation on three datasets. The Merge method is a simple strategy we use to verify the collaboration of internal and external knowledge. It directly generates a passage and merges it into the retrieved passages as the context of the LLM.
@@
Specifically, as the knowledge in LLM’s parameter becomes richer and the abilitiy of the LLM becomes stronger, sometimes answers with search information are not as good as direct answers. To validate the necessity of collaboratively synergizing internal (or parametric) and external (or retrieved) knowledge, we conducted experiments to compare performance. As shown in Fig. 1, across the three evaluation tasks, direct generation and GenRead (Yu et al., 2022) (explicitly generated content) sometimes shows stronger performance. Also, we conduct a test experiment,“Merge”, that explicitly integrates internal and external knowledge by combining retrieved passages with internally generated passages as the final context. “Merge” often achieves better results than both direct generation and RAG approaches, demonstrating the potential of internal and external knowledge collaboration. However, its improvements are not consistent across all datasets, indicating the need for more sophisticated integration methods.
@@
Our contributions are summarized as follows:
@@
2 Related Works
2.1 Retrieval-augmented Generation
In recent years, to address outdated knowledge and hallucination of LLM, RAG has been introduced (Fan et al., 2024; Gao et al., 2023), and many efforts have been made in two aspects: “how to retrieve more relevant information” including retriever fine-tuning (Nian et al., 2024) and query optimization (Ma et al., 2023; Wang et al., 2023a, 2024b) and “how to better use the retrieved information” including domain fine-tuning (Wang et al., 2024c; Zhang et al., 2024; Yue et al., 2025; Xia et al., 2025) and controlled decoding strategies (Shi et al., 2023). Our CoCoA falls into the second category: better utilization of knowledge.
2.2 RAG Pipeline Optimization
Pipeline optimization usually adds pre-generation processing, retrieval intent identification, or optimizes the pipeline as a whole. For example, Glass et al. (2022); Kim and Lee (2024) and Yu et al. (2023) introduce reranking and refinement steps before generation, mitigating the impact of noisy retrieved passages. SKR (Wang et al., 2023b) and UAR (Cheng et al., 2024) avoid unnecessary retrieval by adding retrieval intent identification processes before generation. SUREKim et al. (2024a) first generates multiple candidate answers and performs conditional summary verification based on the candidate answers, allowing LLMs to focus on specific contexts. AstuteRAG (Wang et al., 2024a) integrates reliable information iteratively. However, these methods fail to effectively combine internal and external knowledge or enhance the LLM’s ability to achieve synergy between them, which can limit performance.
2.3 RALM Enhancement
Retrieved-Augmented Language Model(RALM) enhancement is usually achieved by adjusting the LLM to achieve effective use of the information. One common approach is to train the LLM itself. For example, RAFT (Zhang et al., 2024) and InstructRAG (Wei et al., 2024) improve robustness to noisy context via noise-resistance training. REAR (Wang et al., 2024c) balances external and internal knowledge by training relevance-guided generation. Self-RAG (Asai et al., 2023) trains LLMs to decide whether to perform retrieval and to improve its self-reflection capabilities. Another approach involves guiding the decoding (Shi et al., 2023; Kim et al., 2024b). For instance, CAD (Shi et al., 2023) enforces trust in retrieved information by contrastive decodingHowever, both approaches tend to underutilize the model’s internal knowledge, which may constrain the quality and informativeness of its responses.
@@ 3.1 Preliminaries
We formalize the standard RAG framework. Given a query $q$ and a corpus $\mathcal{D}$, the RAG system retrieves $k$ relevant passages $C=\{c_{1},c_{2},\cdots,c_{k}\}\subset\mathcal{D}$ and generates an answer $\hat{a}$ based on the combined input. This process follows a retrieve-then-generate paradigm and can be formulated as:
$\begin{split}C&=\mathcal{R}(q,\mathcal{D},k),\\ \hat{a}&=G(\mathcal{P}(q,C)),\end{split}$ (1)
where $\mathcal{R}$ is the retriever, $\mathcal{P}$ is the prompt constructor that formats $q$ and $C$, and $G$ is the generator (e.g., a LLM) that predicts the final answer $\hat{a}$.
3.2 Two-stage Framework: CoCoA-zero
In this section, we present our multi-agent RAG framework, CoCoA-zero, which also functions as the data synthesis pipeline for CoCoA. As shown in Fig. 2, Stage I (§ 3.2.1) employs two specialized agents to induce knowledge from parameters and retrieval, while Stage II (§ 3.2.2) introduces a agent to synthesize them for high-level decision-making.
3.2.1 Stage I: Knowledge Induction.
It is challenging to extract implicit knowledge solely from the model’s internal knowledge or retrieved passages. Inspired by GenRead (Yu et al., 2022) and SURE (Kim et al., 2024a), we design two dedicated agents for knowledge induction. Each agent first generates an answer and then summarizes knowledge based on that answer.
Internal Knowledge Induction Agent.
Directly allowing the model to explicitly generate its own internal knowledge is difficult to control and will inevitably result in sparse or inconsistent knowledge being generated. Following SURE (Kim et al., 2024a), we introduce conditional induction. Specifically, the Internal Knowledge Agent samples a candidate $a_{\text{in}}$ from the LLM based on the question:
@@ 3.2 Two-stage Framework: CoCoA-zero
External Knowledge Induction Agent.
For retrieved passages, the External Knowledge Agent follows a similar procedure. Specially, it first retrieve some passages $C=\{c_{1},c_{2},\cdots,c_{k}\}$ from the corpus $\mathcal{D}$. Conditioned on both $q$ and $C$, it produces a second candidate $a_{\text{ex}}$:
@@
The conditional knowledge induction framework thus: (1) renders implicit knowledge explicit and controllable; (2) serves as a secondary verification of responses; and (3) establishes a solid foundation for high-level decision-making in the next stage.
3.2.2 Stage II: High-level Decision Making.
@@
Decision-Making Agent.
The Decision-Making Agent adopts COT (Wei et al., 2022) reasoning over the internal and external candidate answers and their corresponding knowledge. It will be prompted with all five components (questions, internal and external candidate answers and their corresponding inductive knowledge) and generate the final answer $\hat{a}$ through COT.
@@ 3.3 Collaborative Chain-of-Agents Training
3.3.1 Supervised Fine-Tuning.
@@
This training explicitly exposes the model to long collaborative samples, where the target outputs are synthesized based on CoCoA-zero. Through end-to-end training, multiple agents can influence and enhance each other’s capabilities. Moreover, the noise introduced by intermediate agents becomes negligible, as it contributes to the overall robustness of the training process.
@@
4 Experiments
In this section, we report our experiments results, and provide a analysis of them. More supplements are in the Appendix.
4.1 Implementation Details
Training Data We sample subsets from the training sets of HotpotQA (Ho et al., 2020a), 2WikiMultiHopQA (Ho et al., 2020b) and WebQuestions (Berant et al., 2013), then synthesize data using the CoCoA-zero and filter them based on gold answers. This results in 6.8k filtered samples for SFT. For DPO, we select 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero framework. For each sample, we gather 5 relevant passages using CONTRIEVER (Izacard et al., 2021).
Training Details We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05). During SFT, we train for 5 epochs with a learning rate of 3e-5. For DPO, we used $\beta$=0.2 and $\alpha$=0.2 (RPO), with a learning rate of 5e-6. All experiments are conducted on a single A100 GPU.
Inference Details During inference, we use Contriever (Izacard et al., 2021) as the retriever and set k to 5. For all datasets, we use 21M English Wikipedia (Karpukhin et al., 2020) dump as the source passages for the retrieval. Prompts for the experiments can be found in Appendix F.
4.2 Datasets and Evaluation Metrics
Evaluation Datasets To evaluate the effectiveness of CoCoA, we conduct experiments on open-domain QA task: WebQuestions (Berant et al., 2013), and TriviaQA (Joshi et al., 2017), as well as multi-hop QA task: HotpotQA (Ho et al., 2020a) and 2WikiMultiHopQA (Ho et al., 2020b). Details are provided in Appendix A.
Evaluation Metrics We report both Exact Match (EM) and F1 scores. Following Asai et al. (2023); Mallen et al. (2022), we adopt a non-strict EM metric that deems a prediction correct if it contains the gold answer. F1 measures token-level overlap between the predicted and gold answers. In our setting, longer responses often yield higher EM scores, but may reduce F1.Thus, considering both metrics provides a more balanced evaluation.
@@ 4.3 Baselines
We selected several of the most representative methods for comparison. (1) StandardRAG, which is the classic “retrieve-then-read” paradigm. (2) Chain-Of-Thought (Wei et al., 2022): Uses CoT prompting to reason before answering. (3) Chain-Of-Note (Yu et al., 2023): Refines the retrieved passages prior to answering. (4) GenRead (Yu et al., 2022): Generates context to answer. (5) SURE (Kim et al., 2024a): Conditional summarization followed by multiple validation. (6) Self-RAG (Asai et al., 2023): Employs adaptive retrieval and self-reflection to decide when and how to use external context. (7) DeepSeek-R1-Distill-8B (Guo et al., 2025): A distilled LLaMA-8B model released by DeepSeek-R1, trained on reasoning data. (8) InstructRAG (Wei et al., 2024): Denoising training using self-synthesized data. All retrieval-based methods use top-5 passages. Details experimental settings are shown in the Appendix C.
@@ 4.4 Main Results
(1) Retrieval vs. non-retrieval.On WikiMQA and WebQuestions, direct generation performs better, while retrieval methods excel on other tasks. This demonstrates that retrieved knowledge and parametric knowledge each have their own strengths and weaknesses in different scenarios.
(2) RAG without training. The improvements of some process optimization methods are decreasing compared to standardRAG. We speculate that this is because current LLMs are becoming more powerful enough to make good use of external knowledge. CoCoA-zero improves the average EM and F1 of all tasks by 4.99% and 4.64% respectively, while other train-free methods show little effect. These results suggest that current QA tasks should place greater emphasis on leveraging the model’s rich internal knowledge.
(3) Superiority and Generalization of CoCoA. Our CoCoA methods achieve state-of-the-art performance across almost all datasets. In particular, CoCoA improves the EM and F1 of 2WikiMultiHopQA tasks by 15.2% and 15.51% respectively. Moreover, despite being trained with limited data, CoCoA also performed well on out-of-distribution dataset, demonstrating its robustness.
(4) Reasoning Distillation vs. CoCoA Training. DeepSeek-R1-8B, trained on distilled reasoning data, outperforms the undistilled StandardRAG. CoCoA, distilled with multi-agent self-synthesis on knowledge-intensive tasks, further surpasses DeepSeek-R1-8B. We speculate this is because logical reasoning and knowledge-intensive tasks differ, and CoCoA can better leverage knowledge. This suggests that explicitly leveraging key internal and external knowledge can be more effective than chain-of-thought reasoning.
(5) Effect of DPO. Comparing our SFT and DPO variants, DPO training yields improvements across several datasets. This suggests that contrastive preference learning can help the model better align to the collaborative responses of multi-agents. However, it may also lead to performance degradation due to the quality of training data.
4.5 Ablation Study I: Different Agents